I am having trouble with this problem.
R varies inversely with S. If R is 40 when S is 2, find S when R is 16.

B) How many digits does the smallest repeating block of the decimal expansion of 5/13 have?
thankyou for your time. Bill
: I am having trouble with this problem.
R varies inversely with S. If R is 40 when S is 2, find S when R is 16.

B) How many digits does the smallest repeating block of the decimal expansion of 5/13 have?
thankyou for your time. Bill

Answer by ankor@dixie-net.com(4489)

R varies inversely with S. If R is 40 when S is 2,
:
You know that inverse means that R decreases when S increases
:
R = k/S
Find k using R=40 & S=2
40 = k/2
40*2 = k; multiply both sides by 2, therefore
k = 80
:
:
find S when R is 16, (we know k=80).
16 = 80/S
16S = 80; multiplied both sides by S
S = 80/16
S = 5
:
B) How many digits does the smallest repeating block of the decimal expansion of 5/13 have?
Just enter 5/13 on good calculator and look: 5/13 = .384615,384615; it would be 6 digits